# Rivers under Additive Noise

### v2

@@ 3 The linear case
so solutions go to +$\infty$+ ∞ (respectively, $\infty$- ∞) exponentially fast with rate $c$ when $x$ is above (respectively below) $\mathscr{R}(0)$ while if $x=\mathscr{R}(0)$,
@@ 4 Main results
where $W$ is a one-dimensional Wiener process defined on a probability space $(\Omega,\mathscr{F},\mathbb{P})$, $s\geq 0$, $\sigma>0$, and $x\in\mathbb{R}$. We will study the long-time behaviour of the unique local solution of (4.1) which we denote by $X_{s,t}(x)$, and we will see that (4.1) possesses both a random repelling river $\mathscr{R}$ (which satisfies $\mathscr{R}(t)-t\to 0$ as in the deterministic case (2.1)) and an attracting river of trajectories which converge to 0 as $t\to\infty$ (again as in (2.1)). Just like in the linear case, $\mathscr{R}(0)$ is random. All trajectories starting above $\mathscr{R}(s)$ at time $s\geq 0$ blow up to $\infty$ in finite time while all trajectories starting below $\mathscr{R}(s)$ converge to 0 as $t\to\infty$. It will turn out that $\mathscr{R}(0)$ will be $\infty$- ∞ with positive probability. In this case, $\mathscr{R}(t)$ will start becoming larger than $\infty$- ∞ at some positive random time and then stay finite for all future times and just like in the nonlinear deterministic and the linear random cases it will solve the underlying equation. All these properties (and more) will follow from four theorems stated below. Further, Proposition 4.10 shows that, unlike the repelling river in the deterministic case in (2.1), $\mathscr{R}(t)$ oscillates around $t$ as $t\to\infty$ almost surely.
@@
where $H:[0,\infty)\to\mathbb{R}$ is a given (random or deterministic) continuous function satisfying $H(0)=0$, $x\in\mathbb{R}$, and $s\geq 0$. Again we denote the unique local solution of (4.2) by $X_{s,t}(x)$.
First of all, arguing from the contrary from (4.2), we have the fundamental monotonicity property:
@@
Note that, for each pair $s<t$, the map $x\mapsto X_{s,t}(x)$ from (,]($\infty$,$\infty$]( - ∞ , ∞ ] to (,]($\infty$,$\infty$]( - ∞ , ∞ ] is continuous.
We define the blow-up time by
@@
Assume that $x\leq 0$. Then
$\inf_{s\leq t\leq v}$ $X_{s,t}(x)\geq\inf_{s\leq t\leq v}\left(x+\int_{s}^{t}X(u)\big{(}X(u)-u\big{)}\,\text{\rm{d}}u\right)+\inf_{s\leq t\leq v}\left(H(t)-H(s)\right)$
$\geq$ $x\wedge 0+\inf_{s\leq t\leq u\leq v}\big{(}H(u)-H(t)\big{)},$
where the inequality
$\inf_{s\leq t\leq v}\left(x+\int_{s}^{t}X(u)\big{(}X(u)-u\big{)}\,\text{\rm{d}}u\right)\geq x\wedge 0$
follows since $y(y-r)\geq 0$ for $r\geq 0$ and $y\leq 0$. The case of $x>0$ follows by monotonicity of Lemma 4.1. ∎
Proposition 4.2 implies that the function $t\mapsto X_{s,t}(x)$ cannot reach $\infty$- ∞ in finite time. Further, if $X_{s,t}(x)\geq A$ for some $t\in[s,v]$ and $A\geq v+\inf_{s\leq\kappa\leq u\leq v}\big{(}H(u)-H(\kappa)\big{)}$, then
@@
To see this, let $H_{*}$ be a lower bound of $H(t)-H(s)$ for $t\in[s,s+1]$ and let $x$ be so large that the solution of $Y^{\prime}=\frac{1}{2}Y^{2}$ with initial condition $x+H_{*}-1$ at time $s$ blows up before time $s+1$. In addition, assume that $x\geq 3-H_{*}$. Let $X(t):=X_{s,t}(x)$, $t\geq s$. Then $X(s)>Y(s)$ since $H_{*}\leq 0$. Let $\tau$ be the infimum of all $t\geq s$ such that $X(t)=Y(t)$ and assume that $\tau\leq s+1$. For $u\leq\tau$ we have $X(u)\geq Y(u)\geq x+H_{*}-1$ since $Y$ is increasing. Then, for $u\in[s,\tau]$,
@@
Theorem 4.5 implies that for each $s\geq 0$ and $x\in\mathbb{R}$ the set $\Omega_{s,x}\subset\Omega$ consisting of those $\omega\in\Omega$ for which the solution starting at $(s,x)$ neither blows up nor converges to zero has measure 0. For $s\geq 0$ denote the union of all $\Omega_{s,x},\,x\in\mathbb{R}$ by $\Omega_{s}$. Since this is an uncountable union it does not follow that $\Omega_{s}$ has measure 0. Indeed we will see in the following that $\mathbb{P}(\Omega_{s})>0$ for every $s\geq 0$. In order to obtain a better understanding of the set $\Omega_{s}$ we will use the notation $I_{1}(s,\omega)$ of those $x\in\mathbb{R}$ for which $\beta_{s}(x)<\infty$ corresponding to a realisation $\omega$, consistent with (4.3). We saw that $I_{1}(s,\omega)$ is a non-empty open interval which is unbounded from above. Similarly, we set
@@
Again, by monotonicity of solutions, $I_{2}(s,\omega)$ is a (possibly empty) interval. Theorem 4.5 implies that every $x\in\mathbb{R}$ is almost surely contained in the (disjoint) union of these two intervals, so the complement of the union of these intervals can contain at most one point. We set
@@
Recall that $\mathscr{R}(s)\in[-\infty,\infty)$ for all $s\geq 0$ and that $\mathscr{R}(s)=-\infty$ implies $\mathscr{R}(r)=-\infty$ for every $r<s$. Further, $X_{s,t}(\mathscr{R}(s))=\mathscr{R}(t)$ whenever $t\geq s$ and $\mathscr{R}(s)>-\infty$, so $\mathscr{R}(t)$ is a solution of (4.1) on the interval where it is larger than $\infty$- ∞. Theorem 4.7 implies that, for each $\varepsilon>0$ and $\alpha\in(0,1/2)$, $\lim_{s\to\infty}\mathbb{P}\big{(}s^{\alpha}|\mathscr{R}(s)-s|\geq\varepsilon\big{)}=0$ but the following result is stronger.
@@
$A$ :=$:=\{\omega:\exists s_{0}\geq 0,\,\forall s\geq s_{0}:\,\mathscr{R}(s)\geq s-\frac{c}{\sqrt{s}}\},$
$B$ :=$:=\{\omega:\exists s_{0}\geq 0,\,\forall s\geq s_{0}:\,\mathscr{R}(s)\leq s+\frac{c}{\sqrt{s}}\},$
@@
and there exists some $T=T(\omega)\in(0,\infty)$ such that $\mathscr{R}_{n}$ converges to $\mathscr{R}$ uniformly on $[T,\infty)$.
@@
such that $Z(t)$ converges to 0 as $t\to\infty$ ($Z$ may be $\infty$- ∞ up to some random time $\zeta$). Then, for $t>\zeta$,
@@
which converges to 0 uniformly on $[T(\omega),\infty)$, so the final assertion follows. ∎
@@
  • The solution with initial condition $(s,x)$ with $|x-s|<1$ will hit one of the two lines parallel to the diagonal at distance 1 for some $t\geq s$. We will provide rather precise asymptotic results for the probability $p^{+}(s,x)$ of hitting the upper line before the lower line in Lemma 5.4. Together with Lemma 5.1, we thus obtain a lower bound for the blow-up probability for these initial conditions namely $B(s,x)\geq p^{+}(s,x)\overline{B}(s)$. These bounds will later turn out also to be asymptotic upper bounds but so far, we can not exclude the possibility that solutions starting at $|x-s|\leq 1$ and hitting the lower set $\{(t,t-1):\,t\geq s\}$ will still return to a neighborhood of the diagonal many times and eventually blow up with large probability (or even almost surely).
@@
  • The solution with initial condition $(s,x)$, $x<0$ will hit 00 after finite time (irrespective of the starting time $s\geq 0$). This is Lemma 5.6.
  • Let $\chi(s)$ be the probability that the solution $X_{s,t}(0)$, $t\geq s$ will remain in $[-1,1]$ for all $t\geq s$ and that $\lim_{t\to\infty}X_{s,t}(0)=0$. Lemma 5.8 states that $\lim_{s\to\infty}\chi(s)=1$. Hence $\overline{\chi}(s):=\inf_{u\geq s}\chi(u)$ converges to 1 as well.
@@
(the $\leq$” is due to the fact that there may be solutions which reach $y$ at time $t$ and which are $\infty$- ∞ at time $s$) converges to 0 as $t\to\infty$, so $\kappa(s,x)=0$ for all $s\geq 0$ and all $x\in\mathbb{R}$ showing that $B(s,x)+C(s,x)\geq 1$ as desired and Theorem 4.5 follows.
@@
Let $b(u):=u(u+s)-1$, and denote by $p$ the scale function of $\overline{Y}$, i.e.
@@
For $x\in[-1,1]$, $Y(t),\,t\geq s$ with initial condition $Y(s)=x$ will exit the interval $[-1,1]$ in finite time, almost surely. Let $p^{+}(s,x)$ be the probability that $Y(t)$, $t\geq s$ exits the interval $[-1,1]$ via 1. Then, for $z\in\mathbb{R}$,
@@
By Lemma A.1a), the probability $p_{A}(s,x)$ that $Z_{A}$ exits the interval $[-1,1]$ via 1 equals
@@
Fix $s>0$ and $z\in\mathbb{R}$ such that $x:=z\frac{\sigma}{\sqrt{2s}}\in[-1,1]$ and consider $Z_{\overline{A}}$ and $Z_{\underline{A}}$ with initial condition $x$ at time $s$. By comparison, we have $Z_{\underline{A}}(t)\leq Y(t)\leq Z_{\overline{A}}(t)$ for $t\in[s,s+s^{\alpha}]$ up to the minimum of the exit times $T(s,x,\overline{A})$ and $T(s,x,\underline{A})$ of $Z_{\overline{A}}$ and $Z_{\underline{A}}$ from $[-1,1]$, where $Y$ denotes the solution of (5.3) with the same initial condition $x$. We saw that, as $s\to\infty$, the probability of exiting the interval $[-1,1]$ via 1 is asymptotically the same for $Z_{\underline{A}}$ and for $Z_{\overline{A}}$. In order to ensure that this is also true for $Y$ we have to show that the exit times from $[-1,1]$ of these processes are at most $s+s^{\alpha}$ with probabilities converging to 1 as $s\to\infty$. We have
$p^{+}(s,x)$ $\geq\mathbb{P}\big{(}Z_{\underline{A}}\mbox{ exits }[-1,1]\mbox{ via }1\mbox{ before or at time }s+s^{\alpha}\big{)}$
$=\mathbb{P}\big{(}Z_{\underline{A}}\mbox{ exits }[-1,1]\mbox{ via}1\big{)}-\mathbb{P}\big{(}Z_{\underline{A}}\mbox{ exits }[-1,1]\mbox{ via }1\mbox{ after time }s+s^{\alpha}\big{)}.$
Define $p^{-}(s,x)$ as the probability that $Y$ starting at $x\in[-1,1]$ at time $s$ exits $[-1,1]$ via -1 (clearly, $p^{+}(s,x)+p^{-}(s,x)\leq 1$ but it is not yet clear that there is equality). Then, analogously,
$p^{-}(s,x)$ $\geq\mathbb{P}\big{(}Z_{\overline{A}}\mbox{ exits }[-1,1]\mbox{ via }-1\mbox{ before or at time }s+s^{\alpha}\big{)}$
$=\mathbb{P}\big{(}Z_{\overline{A}}\mbox{ exits }[-1,1]\mbox{ via }-1\big{)}-\mathbb{P}\big{(}Z_{\overline{A}}\mbox{ exits }[-1,1]\mbox{ via }-1\mbox{ after time }s+s^{\alpha}\big{)}.$
@@
It should not come as a surprise that the upper bound for the expected exit time computed in the previous proof blows up as $\sigma\to 0$. After all, in the deterministic case, for $s$ sufficiently large, there exists an initial condition $(s,x)$ with $x\in[s-1,s+1]$ for which the solution stays in the interval $(s-1,s+1)$ forever.
The two previous lemmas together provide an asymptotic lower bound for the blow-up probability for initial conditions of distance at most 1 from the diagonal: Lemma 5.4 provides an asymptotic bound for the probability of reaching $s+1$ before reaching $s-1$ and Lemma 5.1 states that once we hit $s+1$ the probability of blow-up converges to 1 as $s\to\infty$.
@@
Fix $s\geq 0$ and let $A\geq 0$. Define the processes $X$ and $X_{A}$ on $[s,\infty)$ as follows:
@@
Let $\overline{\tau}:=\inf\big{\{}t\geq s:\,X(t)\geq 1\big{\}}$, $\underline{\tau}:=\inf\big{\{}t\geq s:\,X(t)\leq-1\big{\}}$, $\overline{\tau}_{A}:=\inf\big{\{}t\geq s:\,X_{A}(t)\geq 1\big{\}}$ and $\underline{\tau}_{A}:=\inf\big{\{}t\geq s:\,X_{A}(t)\leq-1\big{\}}$. Note that $X(t)\leq X_{1}(t)$ for $t\in[s,\overline{\tau}_{1}\wedge\underline{\tau}_{1}]$ and $X(t)\geq X_{0}(t)$ and $X_{1}(t)\geq X_{0}(t)$ for $t\geq s$. Therefore, it suffices to show that
@@
Clearly, $\limsup_{t\to\infty}X_{A}(t)\leq 0$ almost surely and $\sup_{t\geq s}X_{A}(t)$ converges to 0 as $s\to\infty$ in probability, so i) and ii) follow taking $A=1$. Observing that the processes $X_{0}$ and $-X_{0}$ have the same law, iii) and iv) follow as well and the proof is complete. ∎
@@
In particular, the probability that $X$ starting at $(s,x)$ with $x<0$ reaches the interval $[0,\infty)$ before or at time $s+2D$ is at least $1/2$ (by Markov’s inequality).
@@
Let us show the upper bound for $\mathscr{R}(s)$. Fix a deterministic positive integer $n$. Let $X_{s}$, $s\in[0,n]$ be the [,)[$\infty$,$\infty$)[ - ∞ , ∞ )-valued solution of (4.1) on $[0,n]$ with final condition $X_{n}=n+n^{-\alpha_{0}}$ (note that $X_{s}$ may take the value $\infty$- ∞ in case the process blows up in finite time in the backward time direction). Define

### v3

@@ 3 The linear case
so solutions go to $\infty$ (respectively, -$\infty$) exponentially fast with rate $c$ when $x$ is above (respectively below) $\mathscr{R}(0)$ while if $x=\mathscr{R}(0)$,
@@ 4 Main results
where $W$ is a one-dimensional Wiener process defined on a probability space $(\Omega,\mathscr{F},\mathbb{P})$, $s\geq 0$, $\sigma>0$, and $x\in\mathbb{R}$. We will study the long-time behaviour of the unique local solution of (4.1) which we denote by $X_{s,t}(x)$, and we will see that (4.1) possesses both a random repelling river $\mathscr{R}$ (which satisfies $\mathscr{R}(t)-t\to 0$ as in the deterministic case (2.1)) and an attracting river of trajectories which converge to 0 as $t\to\infty$ (again as in (2.1)). Just like in the linear case, $\mathscr{R}(0)$ is random. All trajectories starting above $\mathscr{R}(s)$ at time $s\geq 0$ blow up to $\infty$ in finite time while all trajectories starting below $\mathscr{R}(s)$ converge to 0 as $t\to\infty$. It will turn out that $\mathscr{R}(0)$ will be -$\infty$ with positive probability. In this case, $\mathscr{R}(t)$ will start becoming larger than -$\infty$ at some positive random time and then stay finite for all future times and just like in the nonlinear deterministic and the linear random cases it will solve the underlying equation. All these properties (and more) will follow from four theorems stated below. Further, Proposition 4.10 shows that, unlike the repelling river in the deterministic case in (2.1), $\mathscr{R}(t)$ oscillates around $t$ as $t\to\infty$ almost surely.
@@
where $H:[0,\infty)\to\mathbb{R}$ is a given (random or deterministic) continuous function satisfying $H(0)=0$, $x\in\mathbb{R}$, and $s\geq 0$. Again we denote the unique local solution of (4.2) by $X_{s,t}(x)$. We define the blow-up time by
Note that the local solution $X_{s,t}(x)$ depends continuously on $(t,x)$. Arguing from the contrary from (4.2), we have the fundamental monotonicity property:
@@
Assume that $x\leq 0$ and that $u\in[s,v]$ is such that $X_{s,u}(x)\leq x$. Let $t=t(u)\in[s,u]$ be such that $X_{s,t}(x)=x$ and $X_{s,r}(x)\leq x$ for all $r\in[t,u]$ (such a $t$ exists by continuity of $r\mapsto X_{s,r}(x)$ and the fact that $X_{s,s}(x)=x$). Then
$X_{s,u}(x)$ $=X_{s,t}(x)+\int_{t}^{u}X_{s,r}(x)\big(X_{s,r}(x)-r\big)\,\text{\rm{d}}r+H(u)-H(t)\geq x+H(u)-H(t)$
$\geq x+\inf_{s\leq\overline{t}\leq\overline{u}\leq v}\big(H(\overline{u})-H(\overline{t})\big),$
since the integrand is at least 0 since $X_{s,r}(x)\leq x\leq 0$ for $r\in[t,u]$. This holds true for all $u$ as above and therefore the assertion holds for $x\leq 0$.
The case of $x>0$ follows by monotonicity from Lemma 4.1. ∎
Proposition 4.2 implies that the function $t\mapsto X_{s,t}(x)$ cannot reach -$\infty$ in finite time. Further, if $X_{s,t}(x)\geq A$ for some $t\in[s,v]$ and $A\geq v+\inf_{s\leq\kappa\leq u\leq v}\big(H(u)-H(\kappa)\big)$, then
@@
To see this, let $H_{*}$ be a lower bound of H(t)H(s)H(t)-H(s) for $t\in[s,s+1]$ and let $x$ be so large that the solution of $Y^{\prime}=\frac{1}{2}Y^{2}$ with initial condition $x+H_{*}-1$ at time $s$ blows up before time $s+1$. In addition, assume that $x\geq 3-H_{*}$. Let $X(t):=X_{s,t}(x)$, $t\geq s$. Then $X(s)>Y(s)$ since $H_{*}\leq 0$. Let $\tau$ be the infimum of all $t\geq s$ such that $X(t)=Y(t)$ and assume that $\tau\leq s+1$. For $u\leq\tau$ we have $X(u)\geq Y(u)\geq x+H_{*}-1$ since $Y$ is increasing. Then, for $u\in[s,\tau]$,
@@
Theorem 4.5 implies that for each $s\geq 0$ and $x\in\mathbb{R}$ the set $\Omega_{s,x}\subset\Omega$ consisting of those $\omega\in\Omega$ for which the solution starting at $(s,x)$ neither blows up nor converges to zero has measure 0. For $s\geq 0$ denote the union of all $\Omega_{s,x},\,x\in\mathbb{R}$ by $\Omega_{s}$. Since this is an uncountable union it does not follow that $\Omega_{s}$ has measure 0. Indeed we will see in the following that $\mathbb{P}(\Omega_{s})>0$ for every $s\geq 0$. In order to obtain a better understanding of the set $\Omega_{s}$ we will use the notation
$I_{1}(s,\omega):=\{x\in\mathbb{R}\,|\beta_{s}(x)<\infty\}$
corresponding to a realisation $\omega$, consistent with (4.3). We saw that $I_{1}(s,\omega)$ is a non-empty open interval which is unbounded from above. Similarly, we set
@@
Again, by monotonicity of solutions, $I_{2}(s,\omega)$ is a (possibly empty) interval lying below $I_{1}(s,\omega)$. Using Theorem 4.5 and a Fubini-type argument we will show that $I_{2}(s,\omega)$ is unbounded below whenever $I_{2}(s,\omega)$ is nonempty, that $I_{1}(s,\omega)$ is unbounded below whenever $I_{2}(s,\omega)$ is empty and that the complement of the (disjoint) union of these intervals can contain at most one point, almost surely.
Let
$\zeta(s,\omega):=\inf I_{2}(s,\omega)\mbox{ if }I_{2}(s,\omega)\neq\emptyset$
and $\zeta(s,\omega):=$-$\infty$ otherwise. If $P\big(\zeta(s,\omega)>-\infty\big)>0$, then, by monotone convergence, there exists some $x\in\mathbb{R}$ such that $\mathbb{P}\big(\zeta(s,\omega)>x\big)>0$. On the set $\{\omega:\zeta(s,\omega)>x\}$ the number $x$ lies below the set $I_{2}(s,\omega)$ and, a fortiori, also below $I_{1}(s,\omega)$, so $x$ is not in the union of the two intervals which contradicts Theorem 4.5, so $\zeta(s,\omega)=$-$\infty$ almost surely, so $I_{2}(s,\omega)$ is unbounded below whenever $I_{2}(s,\omega)$ is nonempty.
Almost exactly the same argument shows that $\inf I_{1}(s,\omega)=$-$\infty$ on the set where $I_{2}(s,\omega)$ is empty. The fact that, almost surely, the complement of the union of $I_{1}(s,\omega)$ and $I_{2}(s,\omega)$ contains at most one point is clear when $I_{2}(s,\omega)$ is empty (and hence $I_{1}(s,\omega)=\mathbb{R}$). Otherwise, assuming that $\sup I_{2}(s,\omega)<\inf I_{1}(s,\omega)$ with strictly positive probability, by Fubini’s theorem, there must exist some (deterministic) $x\in\mathbb{R}$ for which $\mathbb{P}\big(x\in\big(\sup I_{2}(s,\omega),\inf I_{1}(s,\omega)\big)>0$ so, in particular, $x$ is not contained in the union of $I_{1}(s,\omega)$ and $I_{2}(s,\omega)$ which again contradicts Theorem 4.5.
We set
@@
Recall that $\mathscr{R}(s)\in[-\infty,\infty)$ for all $s\geq 0$ and that $\mathscr{R}(s)=$-$\infty$ implies $\mathscr{R}(r)=$-$\infty$ for every $r<s$. Further, $X_{s,t}(\mathscr{R}(s))=\mathscr{R}(t)$ whenever $t\geq s$ and $\mathscr{R}(s)>$-$\infty$, so $\mathscr{R}(t)$ is a solution of (4.1) on the interval where it is larger than -$\infty$. Theorem 4.7 implies that, for each $\varepsilon>0$ and $\alpha\in(0,1/2)$, $\lim_{s\to\infty}\mathbb{P}\big(s^{\alpha}|\mathscr{R}(s)-s|\geq\varepsilon\big)=0$ but the following result is stronger.
@@
$A$ $:=\{\omega:\exists s_{0}\geq 0,\,\forall s\geq s_{0}:\,\mathscr{R}(s)\geq s-\frac{c}{\sqrt{s}}\},$
$B$ $:=\{\omega:\exists s_{0}\geq 0,\,\forall s\geq s_{0}:\,\mathscr{R}(s)\leq s+\frac{c}{\sqrt{s}}\},$
@@
and there exists some $T=T(\omega)\in(0,\infty)$ such that $\mathscr{R}_{n}$ converges to $\mathscr{R}$ uniformly on $T,\infty)$.
@@
such that $Z(t)$ converges to 0 as $t\to\infty$ ($Z$ may be -$\infty$ up to some random time $\zeta$). Then, for $t>\zeta$,
@@
which converges to 0 uniformly on $T(\omega),\infty)$, so the final assertion follows. ∎
@@
  • The solution with initial condition $(s,x)$ with |xs|<1|x-s|<1 will hit one of the two lines parallel to the diagonal at distance 1 for some $t\geq s$. We will provide rather precise asymptotic results for the probability $p^{+}(s,x)$ of hitting the upper line before the lower line in Lemma 5.4. Together with Lemma 5.1, we thus obtain a lower bound for the blow-up probability for these initial conditions namely $B(s,x)\geq p^{+}(s,x)\overline{B}(s)$. These bounds will later turn out also to be asymptotic upper bounds but so far, we can not exclude the possibility that solutions starting at $|x-s|\leq 1$ and hitting the lower set $\{(t,t-1):\,t\geq s\}$ will still return to a neighborhood of the diagonal many times and eventually blow up with large probability (or even almost surely).
@@
  • The solution with initial condition $(s,x)$, $x<0$ will hit 0 after finite time (irrespective of the starting time $s\geq 0$). This is Lemma 5.6.
  • Let $\chi(s)$ be the probability that the solution $X_{s,t}(0)$, $t\geq s$ will remain in [1,1][-1,1] for all $t\geq s$ and that $\lim_{t\to\infty}X_{s,t}(0)=0$. Lemma 5.8 states that $\lim_{s\to\infty}\chi(s)=1$. Hence $\overline{\chi}(s):=\inf_{u\geq s}\chi(u)$ converges to 1 as well.
@@
(the $\leq$” is due to the fact that there may be solutions which reach $y$ at time $t$ and which are -$\infty$ at time $s$) converges to 0 as $t\to\infty$, so $\kappa(s,x)=0$ for all $s\geq 0$ and all $x\in\mathbb{R}$ showing that $B(s,x)+C(s,x)\geq 1$ as desired and Theorem 4.5 follows.
@@
Let b(u):=u(u+s)1b(u):=u(u+s)-1, and denote by $p$ the scale function of $\overline{Y}$, i.e.
@@
For $x\in[-1,1]$, $Y(t),\,t\geq s$ with initial condition $Y(s)=x$ will exit the interval [1,1][-1,1] in finite time, almost surely. Let $p^{+}(s,x)$ be the probability that $Y(t)$, $t\geq s$ exits the interval [1,1][-1,1] via 1. Then, for $z\in\mathbb{R}$,
@@
By Lemma A.1a), the probability $p_{A}(s,x)$ that $Z_{A}$ exits the interval [1,1][-1,1] via 1 equals
@@
Fix $s>0$ and $z\in\mathbb{R}$ such that $x:=z\frac{\sigma}{\sqrt{2s}}\in[-1,1]$ and consider $Z_{\overline{A}}$ and $Z_{\underline{A}}$ with initial condition $x$ at time $s$. By comparison, we have $Z_{\underline{A}}(t)\leq Y(t)\leq Z_{\overline{A}}(t)$ for $t\in[s,s+s^{\alpha}]$ up to the minimum of the exit times $T(s,x,\overline{A})$ and $T(s,x,\underline{A})$ of $Z_{\overline{A}}$ and $Z_{\underline{A}}$ from [1,1][-1,1], where $Y$ denotes the solution of (5.3) with the same initial condition $x$. We saw that, as $s\to\infty$, the probability of exiting the interval [1,1][-1,1] via 1 is asymptotically the same for $Z_{\underline{A}}$ and for $Z_{\overline{A}}$. In order to ensure that this is also true for $Y$ we have to show that the exit times from [1,1][-1,1] of these processes are at most $s+s^{\alpha}$ with probabilities converging to 1 as $s\to\infty$. We have
[…]
Define $p^{-}(s,x)$ as the probability that $Y$ starting at $x\in[-1,1]$ at time $s$ exits [1,1][-1,1] via -1 (clearly, $p^{+}(s,x)+p^{-}(s,x)\leq 1$ but it is not yet clear that there is equality). Then, analogously,
[…]
@@
It should not come as a surprise that the upper bound for the expected exit time computed in the previous proof blows up as $\sigma\to 0$. After all, in the deterministic case, for $s$ sufficiently large, there exists an initial condition $(s,x)$ with $x\in[s-1,s+1]$ for which the solution stays in the interval (s1,s+1)(s-1,s+1) forever.
The two previous lemmas together provide an asymptotic lower bound for the blow-up probability for initial conditions of distance at most 1 from the diagonal: Lemma 5.4 provides an asymptotic bound for the probability of reaching $s+1$ before reaching s1s-1 and Lemma 5.1 states that once we hit $s+1$ the probability of blow-up converges to 1 as $s\to\infty$.
@@
Fix $s\geq 0$ and let $A\geq 0$. Define the processes $X$ and $X_{A}$ on $s,\infty)$ as follows:
@@
Let $\overline{\tau}:=\inf\big\{t\geq s:\,X(t)\geq 1\big\}$, $\underline{\tau}:=\inf\big\{t\geq s:\,X(t)\leq-1\big\}$, $\overline{\tau}_{A}:=\inf\big\{t\geq s:\,X_{A}(t)\geq 1\big\}$ and $\underline{\tau}_{A}:=\inf\big\{t\geq s:\,X_{A}(t)\leq-1\big\}$. Note that $X(t)\leq X_{1}(t)$ for $t\in[s,\overline{\tau}_{1}\wedge\underline{\tau}_{1}]$ and $X(t)\geq X_{0}(t)$ and $X_{1}(t)\geq X_{0}(t)$ for $t\geq s$ (here, $X_{0}(t)$ and $X_{1}(t)$ correspond to $X_{A}(t)$ with $A=0$ and $A=1$, respectively). Therefore, it suffices to show that
@@
Clearly, $\limsup_{t\to\infty}X_{A}(t)\leq 0$ almost surely and $\sup_{t\geq s}X_{A}(t)$ converges to 0 as $s\to\infty$ in probability, so i) and ii) follow taking $A=1$. Observing that the processes $X_{0}$ and $X_{0}$ have the same law, iii) and iv) follow as well and the proof is complete. ∎
@@
In particular, the probability that $X$ starting at $(s,x)$ with $x<0$ reaches the interval $0,\infty)$ before or at time $s+2D$ is at least $1/2$ (by Markov’s inequality).
@@
Let us show the upper bound for $\mathscr{R}(s)$. Fix a deterministic positive integer $n$. Let $X_{s}$, $s\in[0,n]$ be the $[-\infty,\infty)$-valued solution of (4.1) on $[0,n]$ with final condition $X_{n}=n+n^{-\alpha_{0}}$ (note that $X_{s}$ may take the value -$\infty$ in case the process blows up in finite time in the backward time direction). Define